OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (2, 3). A. x = 2 B. x = 3 C. y = 2 D. y = 3

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

y=3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. A. y = -3x - 13 B. y = -3x + 11 C. y = -3x + 13 D. y = -3x + 1

OpenStudy (anonymous):

@KamiBug @iambatman

OpenStudy (shaadykhamlichi):

D

OpenStudy (kamibug):

D is correct. ^.^

OpenStudy (anonymous):

lol thank you

OpenStudy (shaadykhamlichi):

*-*